OpenStudy (anonymous):

Consider the graph of Bx + Cy = A where A, B, and C are all positive constants. (a) Find the coordinates of the x-intercept. Use upper case letters. Find the coordinates of the y-intercept. Use upper case letters. (b) Is the graph increasing or decreasing?

OpenStudy (anonymous):

on X axis you know Y co ordinates are 0 thus x=(A/B). on Y axis X is 0 thus y=(A/C) graph cant b said whether it's increasing or decreasing until a, b, c are properly mentioned

OpenStudy (anonymous):

i got the answer already but thank you :)